use the confficient to define w/c system hand one sol'vn infinitely many or no solution

1) 2x+y=10
2x-3y=5

2) 4y-6x=1
2y-3x=4
3) x+y=2
2x+2y=4
4) 4x+5y=12
2x-y=6
5) y=3x-5
-6x+2y=-10

Answer by ewatrrr(24785)

 
Hi,
3) x+ y=2
2x+2y=4
5) y=3x-5
-3x +y =-5
-6x+2y=-10 | 3) and 5) The systems are identical Lines, Infinite amount of Solutions
2) 4y-6x=1
2y-3x=4 | multiplying 2nd thru by -2, 0 ≠ -7 NO solution
1) 2x+y=10
2x-3y=5 (4.375,1.25)
4) 4x+5y=12
2x-y=6 (3,0)